Assuming that the average annual return of the three-month T-bills is 3% and the average annual return for the SP500 is 9%. If you have the following investment transactions:

(-10000, 11000)

(-1000, 1200)

a. Which project would you execute if the future cash flow is a sure thing?

b. Discuss and present a possible solution: If the future cash flows are simply expected ones and you consider that the risk of these investments can be comparable to an investment in stock markets, would your decision of part a. change? Why?
Answer #1

Answer:-

Here assuming the investments with more return than the market and taking the discount rate of 12%
R = 3% + 9% = 12%

a)

Given the future cash flow is sure
Investment #1

Year 0 CF = -10000
Year 1 CF = 11000
NPV = -10000 + 11000 / 1.12
NPV = -10000 + 9821.43
NPV = - 178.57
The NPV of this transaction is negative

Investment #2
Year 0 CF = - 1000
Year 1 CF = 1200
NPV = -1000 + 1200 / 1.12
NPV = -1000 + 1071.43
NPV = 71.43
The NPV of this transaction is positive

As the future cash flows are certain the project # 1 future cash flows will be more and one should prefer the project #1 as it will be more profitable than the project #2 which will have less cash flows considering the scale of the project.

b)

If the future cash flows are simply expected ones and you consider that the risk of these investments can be comparable to an investment in stock markets
We will consider the market return of 9% of stock markets which is the S&P return.

investment #1
NPV = -10000 + 11000 / 1.09
NPV = 91.74
Investment #2
NPV = -1000 + 1200 / 1.09
NPV = 100.92


As we have to use the market returns which is 9 % the investment decision will change as the Project #2 will be preferred as it has high NPV than project#1 ( 100.92 > 91.74) and it is a one year investment ( return expected ones).